OpenStudy (anonymous):

Express the following as a function of its correlated acute angle and evaluate. tan(39pi/4)

zepdrix (zepdrix):

sammeeeyyyy, what's up? stuck on this?

zepdrix (zepdrix):

So you have 39 of these pi/4's ? That's too many, gotta unwind that angle.

zepdrix (zepdrix):

let's spin backwards one full rotation by subtracting 2pi,\[\Large\rm \frac{39\pi}{4}-2\pi=\frac{39\pi}{4}-\frac{8\pi}{4}=\frac{31\pi}{4}\]Mmm that's no good, we gotta spin back a few more times, ya?

OpenStudy (anonymous):

I got 7pi/4, after unwind it again and again... @zepdrix

zepdrix (zepdrix):

Ah yes that looks right

OpenStudy (anonymous):

What should I do next?

zepdrix (zepdrix):

\[\Large\rm \tan\left(\frac{7\pi}{4}\right)\]AHHH I dunno, this is just unit circle stuff... lot of memorization...

zepdrix (zepdrix):

All of the pi/4's give us a 1 or a -1 when we take the tangent of them.

zepdrix (zepdrix):

We're in which quadrant with our angle?

OpenStudy (anonymous):

Quadrant 4?

zepdrix (zepdrix):

Mmm k good. If you can't seem to get the sign right, you can relate tangent back to sine and cosine. \[\Large\rm \tan=\frac{\sin}{\cos}\]What's the sign going to be in quadrant 4? we getting a 1 or a -1?

OpenStudy (anonymous):

-1

zepdrix (zepdrix):

yay good job \c:/